OpenStudy (anonymous):

assuming x > 0 what is the simplified fraction form for 5x sqrt 3/ sqrt 48x^2?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

\[\frac{ 5x \sqrt{3} }{ \sqrt{48} x^2 }\] Is this what you mean?

OpenStudy (anonymous):

yea

OpenStudy (luigi0210):

First we need to simplify the bottom radical, can you think of any factors of 48?

OpenStudy (anonymous):

2

OpenStudy (luigi0210):

any factors that involve a 3?

OpenStudy (luigi0210):